A runner hopes to complete the 10,000-m run in less than 30.0 min. After running at constant speed for exactly 27.0 min , there are still 1100m to go.

The runner must then accelerate at 0.23 m/s2 for how many seconds in order to achieve the desired time?
Answer #1

TOTAL distance = 10,000 m

initial time = 27 min

in this time speed is constant. so acceleration is zero , S = V0 * t + 1/2 * a * t^2

                        S = V0 * t + 0

initial speed = distance / time = 10,000 - 1100 / 27 * 60

                                   V0 = 8900 / 1620 = 5.49 m/sec

speed of runner to compleate 1100 m in 3 min

                     V = 1100 / 3*60 = 1100 / 180 = 6.11m/sec

acceleration = a = 0.23 m/sec

V = V0 + a t

time = V - V0 / a   = 6.11 - 5.49 / 0.23 = 2.69 sec = 2.7 sec